# Maxime Maria 1, Valentin Guillaume 2, Simon Guionnière 2, Nicolas Dacquay 2, Cyprien Plateau–Holleville 1, Vincent Larroque 1,3, Jean Lardé 2,3, Yassine Naimi 3, Jean-Philip Piquemal 4,5,6, Guillaume Levieux 7, Nathalie Lagarde 2, Stéphane Mérillou 1 and Matthieu Montes 2,6,*

### v1

@@ Introduction
Recent advances in the determination of atomic resolution molecular structures and assemblies using CryoEM [1], in protein structure prediction methods like AlphaFold [2, 3], and the growing availability of molecular dynamics simulations [4, 5], have significantly increased the volume of molecular structural biology data. Additionally, modern high-performance computing hardware has enabled the simulation of larger molecular systems [6, 7, 8], further amplifying data production. This massive molecular data poses a challenge for existing software solutions in terms of data processing, visualization, and storage.
@@ Methods
We compared the performance of VTX version 0.4.4 [9], VMD version 1.9.3 [20], PyMOL version 3.1 [21] and ChimeraX version 1.9 [22] in terms of system file loading time and frame rate using the 2023 Martini minimal whole cell model [7]. This system is a coarse-grained model of 114 million Martini beads available in the Gromacs [23] .gro file format that pre-figures future massive molecular systems constructed for molecular dynamics simulation. The system, displayed in figure 3 contains 60,887 soluble proteins, 2,200 membrane proteins, 503 ribosomes, a single 500 kbp circular dsDNA, 1.3 million lipids, 1.7 million metabolites and 14 million ions. For clarity of the visual analysis, we did not display the 447 million water beads.
The benchmark results are presented in table 1. ChimeraX and PyMOL were unable to load the system due respectively to a crash and a freeze. VMD and VTX loaded the system with similar times and successfully displayed it, but VTX achieved significantly higher frame rate. Any attempts to modify VMD rendering settings caused in a severe drop in frame rate, eventually leading to a complete freeze. With a stable interactive frame rate and despite limited computational resource, VTX allowed to interactively manipulate the system, perform precise selections and modify rendering settings. A video displaying a free-fly real-time navigation inside this system is available as a supplementary material.
@@
Figure 3: Wide-system view of the 2023 Martini minimal whole cell model rendered in real-time with VMD on the left, and VTX on the right. These pictures were captured with the graphical settings used to perform the frame rate measurement in Table 1.
@@ Conclusion
All .gro files that constitute the 2023 Martini minimal whole cell model are available at https://github.com/marrink-lab/Martini_Minimal_Cell.

### v2

@@ Introduction
Recent advances in the determination of atomic resolution molecular structures and assemblies using CryoEM [1], in protein structure prediction methods like AlphaFold [2, 3], and the growing availability of molecular dynamics simulations [4, 5], have significantly increased the volume of molecular structural biology data. Additionally, modern high-performance computing hardware has enabled the simulation of larger molecular systems [6, 7, 8, 9], further amplifying data production. This massive molecular data poses a challenge for existing software solutions in terms of data processing, visualization, and storage.
@@ Methods
We compared the performance of VTX version 0.4.4 [10], VMD version 1.9.3 [21], PyMOL version 3.1 [22] and ChimeraX version 1.9 [23] in terms of system file loading time and frame rate using the 2023 Martini minimal whole cell model [7]. This system is a coarse-grained model of 114 million Martini beads available in the Gromacs [24] .gro file format that pre-figures future massive molecular systems constructed for molecular dynamics simulation. The system, displayed in figure LABEL:fig:rendering contains 60,887 soluble proteins, 2,200 membrane proteins, 503 ribosomes, a single 500 kbp circular dsDNA, 1.3 million lipids, 1.7 million metabolites and 14 million ions. For clarity of the visual analysis, we did not display the 447 million water beads.
The benchmark results are presented in table 1. ChimeraX and PyMOL were unable to load the system due respectively to a crash and a freeze. VMD and VTX loaded the system with similar times and successfully displayed it, but VTX achieved significantly higher frame rate. Any attempts to modify VMD rendering settings caused in a severe drop in frame rate, eventually leading to a complete freeze. With a stable interactive frame rate and despite limited computational resource, VTX allowed to interactively manipulate the system, perform precise selections and modify rendering settings.
A video displaying a free-fly real-time navigation inside this system is available as a supplementary material.
@@
Figure 3: Wide-system view of the lipids, membrane proteins, and ribosomes (36,566,468 Martini beads) of the 2023 Martini minimal whole cell model rendered in real-time with VMD on the left, and VTX on the right. These illustrations were captured on a Dell precision 5480 laptop equipped with an Intel i7-13800H, 32GB of RAM and a NVidia Quadro RTX 3000 GPU. All attempts to improve the rendering with VMD on this setup resulted in a freeze of the application
@@
VMD, ChimeraX and PyMOL are reference molecular visualization tools that are widely used by the community and that, compared to VTX, offer many additional functionalities notably through the use of numerous and diverse plugins developed over time. In the next version of VTX, Python bindings will be supported to allow the development and integration of external scripts and plugins.
@@ Conclusion
All .gro files that constitute the Martini minimal whole cell model are available upon request at the Marrink lab. The procedure to generate the model is available at https://github.com/marrink-lab/Martini_Minimal_Cell.